### v2

# Prediction Intervals for Interim Events in Randomized Clinical Trials with Time-to-Event Endpoints

Abstract
Time-to-event endpoints are central to evaluate treatment efficacy across many disease areas. Many trial protocols include interim analyses within group-sequential designs that control type I error via spending functions or boundary methods, with operating characteristics determined by the number of looks and the information accrued. Planning interim analyses with time-to-event endpoints is challenging because statistical information depends on the number of observed events, so adequate follow-up to accrue the required events is critical and interim prediction of information at scheduled looks and at the final analysis becomes essential. While several methods have been developed to predict the calendar time required to reach a target number of events, to the best of our knowledge there is no established framework that addresses the prediction of the number of events at a future date with corresponding prediction intervals. Starting from prediction interval approach originally developed in reliability engineering for the number of future component failures, we reformulated and extended it to the context of interim monitoring in clinical trials. This adaptation yields a general framework for event-count prediction intervals in the clinical setting, taking the patient as the unit of analysis and accommodating a range of parametric survival models, patient-level covariates, stagged entry and possible dependence between entry dates and loss to follow-up. Prediction intervals are obtained in a frequentist framework from a bootstrap estimator of the conditional distribution of future events. The performance of the proposed approach is investigated via simulation studies and illustrated by analyzing a real-world phase III trial in childhood acute lymphoblastic leukaemia.
00footnotetext: $\dagger$ Correspondence: e.ratti9@campus.unimib.it
Keywords Interim Analysis, Event-driven trials, Prediction Interval, Parametric Bootstrap, Survival Extrapolation
1 Introduction
Time-to-event endpoints are often used for evaluating treatment efficacy in many disease areas. However, designing trials in which the primary endpoint is a time-to-event variable is challenging. In this context, the statistical information of interest depends on the number of events that occur rather than on the number of patients enrolled. Accordingly, the trial is usually powered on the expected number of events [80]. Interim analyses are increasingly planned to allow early stopping for efficacy or futility for ethical and economic reasons [77]. Group-sequential designs are typically adopted and type-I error is controlled via alpha-spending or boundary methods [42]. The operating characteristics of a group sequential test - power, error rates, and stopping probabilities - depend on the number and timing of looks, on the information accrued at each analysis and on the maximum planned information at the final analysis. Consequently, the protocol must ensure sufficient follow‐up to achieve the target number of events required to preserve the nominal power. In addition, interim analyses should be conducted at the information levels specified at the design stage. If analyses are performed at times corresponding to a different number of events than planned, the Type I error probability may be adversely affected [46] [60].
However, achieving both aims is difficult in practice. Sample size calculation relies on the correct specification of the survival distribution, yet trustworthy informations on survival estimates are often unavailable at the design stage and may shift as data mature. When the design assumptions prove inaccurate while the trial is ongoing, the operational characteristics of the study could be jeopardize. Since information accrues when events occur, an unforeseen slowdown in event rates can shift interim analyses to information times that depart from the protocol plan or push completion well beyond the planned timeline. This has an impact on the budget and logistics of the trial, even when the nominal power of the trial is preserved. On the other hand, an unexpected increase in event rates may put the trial at risk of being over-run [78] and appropriate methods should be considered [7]. A disciplined interim monitoring framework can identify such discrepancies early and support timely corrective actions.
In light of these challenges, it is important to prospectively predict while the trial is ongoing the information expected at each planned interim analysis, the maximum (final) information at the scheduled final analysis or the calendar time at which these targets will be reached. If a maximum-information approach is adopted, the trial ends once the target number of events has accrued [44]. A natural predictand is therefore the calendar time required to reach that target, often termed milestone or time to endpoint maturation [57] [76]. In a maximum-duration approach, follow-up is limited by a fixed calendar time [44]. In this case, natural predictands are the cumulative number of events or, equivalently, the additional number of events expected by future calendar dates and at the planned study end.
The point and interval prediction of the time to endpoint maturation has been extensively studied in the literature [6, 83, 81, 82, 66, 16, 47, 52]. Heitjan et al. [34] provide a comprehensive review of existing methods, highlighting the importance of prediction intervals for time to endpoint maturation to quantify uncertainty around the study’s closing date. Some approaches accommodate ongoing accrual at the time of prediction [76, 85, 3].
Predicting the additional number of events at interim time points is critical for decision-making, especially when new drugs are being developed at a fast pace and adaptive and complex trials designs are increasingly needed, e.g. platform trials. Neverthells, most available methods in the biostatistical literature provide only point prediction under a few survival distributions. Fang and Su [24] proposed a hybrid approach that combines nonparametric and parametric modeling to estimate the survival curve. Specifically, change points in the survival function are first detected using a data-driven algorithm developed by Goodman et al. [28]. The segment preceding the final change point is estimated nonparametrically, while the tail beyond it is modeled using a parametric exponential or Weibull distribution. This approach is implemented by Rufibach [63] in the eventTrack R package. A growing body of literature has also addressed the use of point prediction for the purpose of sample size re-estimation [79, 31, 72, 25, 56]. Even though point predictions are easy to communicate they do not quantify uncertainty about reaching the planned information level and target number of events by the prespecified date. Prediction intervals address this limitation by explicitly reflecting uncertainty through their limits and width, making them more informative for interim trial monitoring. To the best of our knowledge, only two approaches address directly the issue of providing prediction interval for the additional number of events. Anisimov [3] derived closed-form prediction intervals for the number of future events in ongoing multicenter time-to-event trials, explicitly accounting for both newly recruited and at-risk patients under varied recruitment scenarios. The approach models center-level accrual via a Poisson–gamma process and event dynamics with continuous-time finite Markov chains under an exponential model. Anisimov et al. [2] extended the framework to incorporate a cure fraction, under both exponential and Weibull assumptions. However, these approaches construct prediction intervals using the so-called plug-in method [55], i.e., they treat the model parameters as if they were known and simply replace them with their point estimates when computing the conditional distribution of the additional number of events. This has been shown to yield intervals whose coverage can be far from the nominal level [71]. Motivated by a recently completed phase III trial in childhood acute lymphoblast leukemia (ALL), we propose here a general framework to construct prediction intervals for the additional number of events at interim analysis in fixed follow-up times clinical trials by adapting and extending a method originally developed in reliability engineering for predicting component failures [71]. We extended the approach to the clinical setting by treating the patient (rather than a component cohort) as the unit of analysis, accommodating possible dependence between entry dates and loss to follow-up and leveraging patient-level covariates available for instance in case of at unblinded interim looks.
This paper is organized as follows: Section 2 introduces the notation, the prediction framework and briefly reviews the methods originally proposed by Tian et al. [71] for obtain prediction intervals; Section 3 outlines our proposed adaptation and extensions; Section 4, presents the results of a simulation study to evaluate the performance of the modified framework; Section 5 include an application to a phase III trial in childhood acute lymphoblast leukemia (ALL); Section 6 discusses the implications of our findings and highlights potential extensions.
2 Background
The framework proposed by Tian et al. [71] was developed in the context of reliability engineering. Given a sample of items inspected at a given time, the aim is to provide prediction intervals for the number of inspected items that will fail within a future time window, which constitutes a within-sample prediction problem. The authors consider two reliability settings: a single-cohort scenario, where a homogeneous set of units enters testing at the same time, and a multiple-cohort scenario, where homogeneous sets enter at different times.
For both scenarios, Tian et al. [71] propose three bootstrap-based approaches that extend the prediction-interval constructions reviewed in Appendix, Section 7.1 to within-sample prediction with censored survival data: the direct bootstrap, building on [33] [65], the calibration bootstrap following [10], and a generalized pivotal-quantity approach, which uses bootstrap to approximate the distribution of a generalized pivotal quantity for $\bm{\vartheta}$ in log–location–scale models, following [32] [74].
In this paper, we adapt and extend the direct bootstrap method in the multiple-chort scenario. Tian et al. [71] shown that the method has stronger empirical performance than the calibration bootstrap. Moreover, unlike the generalized pivotal-quantity bootstrap—whose construction is primarily well-defined under Type II censoring and only approximated under Type I censoring—the direct bootstrap can accommodates more general censoring mechanisms. We therefore focus on this approach and briefly describe it below.
2.1 Prediction Intervals Definition
Let $\bm{D}$ denote the observed sample of $n$ independent and identically distributed observations arising from a time-to-event process and, possibly, an additional censoring process, with joint density indexed by the parameter vector $\bm{\vartheta}$. Here, $\bm{\vartheta}$ collects the parameters governing the time-to-event distribution and, when applicable, those of the censoring distribution. At a fixed time $t_{c}$ we are interested in predict the value of a future discrete random variable $Y$,i.e. the predictand, here the additional number of events, which has conditional cumulative mass function (CMF) $F_{Y}(y|\boldsymbol{D};\bm{\pi})$, depending on $\bm{\pi}=\varphi(\bm{\vartheta})$. We introduce the notation $\mathscr{D}$, $L(\mathscr{D})$ and $U(\mathscr{D})$ for the sample space of $\boldsymbol{D}$, the lower and upper prediction bounds for $Y$ respectively. If $L(\mathscr{D})$ and $U(\mathscr{D})$ are determined so that
$P(L(\mathscr{D})\leq Y\leq U(\mathscr{D}))=1-\alpha,$ (1)
then $\mathrm{PI}_{1-\alpha}(\mathscr{D})=\ [L(\mathscr{D})\ ,U(\mathscr{D})\ ]$ is called a $1-\alpha$ level prediction interval for $Y$ and $\mathrm{PI}_{1-\alpha}(\bm{D})$ denote its realization. We restrict our attention to two-sided prediction intervals, thus requiring that $L(\mathscr{D})$ and $U(\mathscr{D})$ be both finite. A prediction interval is exact if (1) is exactly equal to $1-\alpha$, while it is asymptotically correct if $\lim_{n\rightarrow\infty}P(L(\mathscr{D})\leq Y\leq U(\mathscr{D}))\rightarrow 1-\alpha$, as is often the case for a discrete predictand.
[…]
2.2 The Direct-bootstrap, Multiple Chort Method (DB-MC)
At the prediction time $t_{c}$, there are $g$ cohorts (groups) of items. Cohort $j$ contains $n_{j}$ items that were put on test at the common known entry time $\tau_{aj}$, with $j=1,\ldots,g$. Let $T_{ij}$ denote the failure time of item $i$ in cohort $j$, and assume
$T_{ij}\ \stackrel{{\scriptstyle\text{iid}}}{{\sim}}\ f_{\boldsymbol{\theta}}(t),\qquad i=1,\ldots,n_{j},\ \ j=1,\ldots,g,$ (2)
where all items share the same (common) PDF $f_{\boldsymbol{\theta}}(t)$. Define the event indicator at $t_{c}$ for item $i$ in cohort $j$ as $\tilde{\delta}_{ij}=\mathbb{I}\!\left(T_{ij}\leq\tau_{j}\right)$ where $\tau_{j}=t_{c}-\tau_{aj}$. The number of observed failures in cohort $j$ is $\delta_{j}=\sum_{i=1}^{n_{j}}\tilde{\delta}_{ij}$ and the number of right-censored items is $\gamma_{j}=n_{j}-\delta_{j}$. The data observed at $t_{c}$ can be represented as a collection of single-cohort datasets, $\bm{D}=\{D_{1},\ldots,D_{g}\}$. For cohort $j$, we summarize the observed information as $D_{j}=(\bm{T}_{j},\delta_{j},\gamma_{j})$, where $\bm{T}_{j}$ is the vector of observed times possibly right-censored, $\delta_{j}$ denotes the number of failures, and $\gamma_{j}$ the number of units censored at $t_{c}$.
Let $Y$ be the number of item failures to be predicted at a future time $\tau_{j}+\Delta t$ on $n=\sum_{j=1}^{g}n_{j}$ tested unit. We denote $\zeta_{j}=\sum_{i=1}^{n_{j}}\mathbb{I}(\tau_{j}<T_{ij}\leq\tau_{j}+\Delta t)$. Since $\zeta_{j}\sim\mathrm{Bin}\left(\pi_{j},\gamma_{j}\right)$ are mutually independent and $\pi_{j}=P(T_{j}\leq\tau_{j}+\Delta t\mid T_{j}>\tau_{j})$, then
$Y=\sum_{j=1}^{g}\zeta_{j}\sim\mathrm{PoiBin}(\bm{\pi},\bm{\gamma}),$ (3)
where $\mathrm{PoiBin}(\bm{\pi},\bm{\gamma})$ denotes the probability mass function (PMF) of a Poisson-binomial distribution with probability vector $\bm{\pi}=(\pi_{1},\ldots,\pi_{g})$, weight vector of the at risk units $\bm{\gamma}=(\gamma_{1},\ldots,\gamma_{g})$ [39] [69]. We briefly review the PMF and the CMF of the Poisson-binomial distribution in the Appendix, Section 7.2.
Following [33], Tian et al. [71] proposed obtaining the $1-\alpha$ prediction interval for $Y$ as the corresponding quantiles of the bootstrap estimator of the conditional CMF. Specifically, once $\bm{D}=\{D_{1},\ldots,D_{g}\}$ is observed, estimate $\bm{\hat{\theta}}$ for $\bm{\theta}$ is found by maximizing the parametric likelihood
$\mathcal{L}(\bm{\theta})$ $\propto\prod_{j=1}^{g}\prod_{i=1}^{n_{j}}f_{\bm{\theta}}(t)^{\tilde{\delta_{i}}}\Big[1-F_{\bm{\theta}}(t)\Big]^{1-\tilde{\delta_{i}}}$, (4)
where $F_{\bm{\theta}}(t)$ is the CDF of $T_{ij}$. Thus, one generates $B$ bootstrap replications of the multiple-cohort data from $f_{\bm{\hat{\theta}}}(t)$ and, for each replication $b=1,\dots,B$, computes the vector $\hat{\boldsymbol{\pi}}^{*}_{b}$ of cohort-specific conditional event-probability at a future prediction horizon $\Delta t$. For each bootstrap sample, one then evaluates the CMF of $Y$, denotes as $F_{Y}(y\mid\hat{\boldsymbol{\pi}}^{*}_{b})$. Averaging these over $b$ yields the direct-bootstrap estimator of the conditional CMF is [71]
$\hat{F}^{*}_{Y}(y\mid\bm{D})=\frac{1}{B}\sum_{b=1}^{B}F_{Y}(y\mid\bm{D}_{b}^{*};\hat{\boldsymbol{\pi}}^{*}_{b}).$ (5)
Thus, a $1-\alpha$ equal-tails prediction interval is given by
$\mathrm{PI}_{1-\alpha}(\bm{D})=\{y:y^{*}_{\alpha/2}\leq y\leq y^{*}_{1-\alpha/2}\},$ (6)
where $\mathscr{D}=\bm{D}$ denotes the observed sample and $y^{*}$ and $y^{*}_{1-\alpha/2}$ are the $\alpha/2$ and $1-\alpha/2$ quantiles of the bootstrap estimator of the CMF of $Y$ (See Appendix, Section 7.1 for details). In terms of coverage, the method is supported by theoretical guarantees of asymptotic nominal coverage. (Theorem 1, 2, 3 in [71]).
3 Extension to the Interim Monitoring Setting
The DB–MC method in Section 2.2 closely matches the clinical trial setting, as it naturally accommodates staggered entry and censoring at the interim cutoff by treating each patient enrolled up to the interim date as a cohort of size one. This combination of patient-specific entry times and a common prespecified cutoff that sets the admissible follow-up is often referred to as generalized type I censoring [45]. In interim monitoring, however, additional right censoring mechanisms may operate beyond generalized type I censoring, most notably random censoring due to loss to follow-up. A further key distinction from the original DB–MC setting is that phase III trials typically exhibit patient-level heterogeneity—driven by randomized treatment allocation and other covariates—which can be leveraged to improve prediction, especially under unblinded monitoring for early efficacy or futility. These extensions are the focus of our work. We consider phase III trials with closed accrual at the prediction time, consistent with our motivating example and Tian et al.[71].
3.1 Methods
At the interim time $t_{c}$, accrual is closed and $n$ patients are available, each with a specific entry time $\mathcal{T}_{aj}$, $j=1,\ldots,n$. Let $T_{j}$ denote the time to event and $L_{j}$ the time to loss to follow-up, both measured from study entry. Define the individual censoring time at the interim as $\mathcal{T}_{j}=t_{c}-\mathcal{T}_{aj}$. Let $C_{j}=\min(L_{j},\mathcal{T}_{j})$, $T_{j}^{\mathrm{obs}}=\min(T_{j},C_{j}),$ and define the event-type indicators $\delta_{j}=\mathbb{I}(T_{j}\leq C_{j})$, $\varepsilon_{j}=\mathbb{I}\big(L_{j}\leq\min(T_{j},\mathcal{T}_{j})\big)$, $\gamma_{j}=\mathbb{I}\big(\mathcal{T}_{j}\leq\min(T_{j},L_{j})\big)$. Finally, let $\bm{z}_{j}$ denote the vector of patient-level covariates. Here we focus on baseline characteristics. We impose the following assumptions on $(T_{j},L_{j},\mathcal{T}_{j})$:
$\text{(A1) }\,\,T_{j}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\ f_{\bm{\theta}}(\,t\mid\bm{z}_{j}).$ (7)
$\text{(A2) }\,\,(L_{j},\mathcal{T}_{j})\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\ g_{\bm{\psi},\bm{\phi}}(l,\tau).$
$\text{(A3) }\,\,T_{j}\perp(L_{j},\mathcal{T}_{j})\ \mid\ \bm{z}_{j}.$
$\text{(A4) }\,\,\bm{\vartheta}=(\bm{\theta},\bm{\psi},\bm{\phi})\in\Theta\times\Psi\times\Phi.$
Note that, since patients enter the study at random calendar times, the distribution of $\mathcal{T}_{j}$ is induced by the distribution of entry times $\mathcal{T}_{aj}$, shifted by $t_{c}$ (i.e., $\mathcal{T}_{j}=t_{c}-\mathcal{T}_{aj}$); see [43]. Thus, we assume an independent right-censoring mechanism, identical across treatment groups and covariate levels and that the PDFs of $T_{j}$, $L_{j}$ and $\mathcal{T}_{j}$ do not share any parameters. In contrast, interim censoring and loss to follow-up are not required to be independent processes, that is, we do not require that $g_{\bm{\psi},\bm{\phi}}(l,\tau)=\tilde{g}_{\bm{\psi}}(l)\,r_{\bm{\phi}}(\tau)$, where $\tilde{g}_{\bm{\psi}}(l)$ and $r_{\bm{\phi}}(\tau)$ are the marginal distribution of $L_{j}$ and $\mathcal{T}_{j}$ respectively. This implies that we do not account for any potential dependence between the loss-to-follow-up times and the entry times, e.g., patients enrolled near the end of accrual may therefore be less likely to be loss to follow-up [43].
Let $Y$ be the number of events to be predicted at a future time $\tau_{j}+\Delta t$ where $\tau_{j}=t_{c}-\tau_{aj}$ denotes a realization of $\mathcal{T}_{j}$. Furthermore, let $\zeta_{j}$ be the indicator function for $T_{j}\leq\mathrm{min}(\tau_{j}+\Delta t,L_{j})$. The probability that a patient will experience the event by the future time $\tau_{j}+\Delta t$, given that neither he/she has not yet experienced the event at the time of the interim $t_{c}$ neither was loss to follow-up is $\pi_{j}=\Pr\bigl(T_{j}\leq\mathrm{min}(\tau_{j}+\Delta t,L_{j})\mid\gamma_{j}=1,\bm{z}_{j}\bigr)$. Under assumptions (7), $\pi_{j}$ can be expressed as
$\pi_{j}=\frac{\displaystyle\big[F_{\boldsymbol{\theta}}(\tau_{j}+\Delta t\mid\mathbf{z}_{j})-F_{\boldsymbol{\theta}}(\tau_{j}\mid\mathbf{z}_{j})\big]-\int_{\tau_{j}}^{\tau_{j}+\Delta t}\tilde{G}_{\boldsymbol{\psi}}(u)\,f_{\boldsymbol{\theta}}(u\mid\mathbf{z}_{j})\,\mathrm{d}u}{\displaystyle\big[1-F_{\boldsymbol{\theta}}(\tau_{j}\mid\mathbf{z}_{j})\big]\,\big[1-\tilde{G}_{\boldsymbol{\psi}}(\tau_{j})\big]},$ (8)
where $F_{\bm{\theta}}$ and $\tilde{G}_{\bm{\psi}}$ are the CDFs of the event time and loss to follow-up time respectively. Proof of (8) and special cases are given in Appendix, Section  7.3. Since $\zeta_{j}\sim\mathrm{Ber}\left(\pi_{j}\right)=\mathrm{Bin}\left(\pi_{j},1\right)$ and are assumed mutually independent, the number of events to be predicted at a future time $\tau_{j}+\Delta t$ has still a Poisson-binomial distribution, that is
$Y=\sum_{j:\gamma_{j}=1}\zeta_{j}\sim\mathrm{PoiBin}(\bm{{\pi}}),$ (9)
were $\bm{{\pi}}$ is the vector of individual conditional probabilities obtained for only the interim censored patients as in (8). Prediction bounds are then obtained as in (6) after having recovered the bootstap estimator of the conditional CMF. Once the interim data $\bm{D}=(T_{j}^{\mathrm{obs}},\delta_{j},\varepsilon_{j},\gamma_{j},\bm{z}_{j})_{j=1}^{n}$ is observed, estimates for $\bm{\theta}$ and $\bm{\psi}$ are found by maximizing the following parametric likelihoods
$\mathcal{L}(\bm{\theta})$ $\propto\prod_{j=1}^{n}f_{\bm{\theta}}(t\mid\bm{z}_{j})^{\delta_{j}}\Big[1-F_{\bm{\theta}}(t\mid\bm{z}_{j})\Big]^{1-\delta_{j}},$ (10)
$\mathcal{L}(\bm{\psi})$ $\propto\prod_{j=1}^{n}\tilde{g}_{\bm{\psi}}(l)^{\varepsilon_{j}}\Big[1-\tilde{G}_{\bm{\psi}}(l)\Big]^{1-\varepsilon_{j}},$
Maximizing the likelihoods in (10) is justified under assumptions (7) and yields the maximum likelihood estimator for $\bm{\theta}$ and $\bm{\psi}$. This justification is given in the Appendix, Section  7.4.
3.2 Proposed Sampling Scheme
The original DB–MC method requires generating many replicates of the multiple-cohort dataset and, for each replicate, computing the vector of conditional probabilities $\bm{\pi}$ at a future prediction horizon $\Delta t$. Translating this step to interim monitoring—where each patient can be viewed as a cohort of size one—requires additional care. We propose a parametric bootstrap sampling scheme to obtain the bootstrap estimator if the conditional CMF of $Y$. The proposed scheme is exposed in the Appendix, Section 7.5.
Some remarks are worth noting regarding the nature of the inference considered in this particular contest. The proposed sampling scheme conditions on the event and censoring indicators, i.e, they are kept fixed across bootstrap replicates. If censoring indicators were different as in classic survival bootstrap [23, 36, 1, 30, 11], the risk set would vary across replicates and the Monte Carlo average in (5) would mix conditional CMFs with different supports. This would target an unconditional CMF, whereas in a within-sample prediction problem it is required the CMF of $Y$ conditional on the observed risk set that defines its support. Conditioning on the indicators also brings information on the observed times. Moreover, each $\pi_{j}$ is conditional on $\tau_{j}$. Hence, the bootstrap must preserve this conditioning in order to recover the sampling distribution of $\boldsymbol{\pi}$ at a future prediction horizon $\Delta t$. Prior observing the interim data $\bm{D}$, $(T_{j}^{\mathrm{obs}},\delta_{j},\varepsilon_{j},\gamma_{j})_{j=1}^{n}$ are random variables with PDF $\mathrm{Pr}_{\bm{\vartheta}}(T_{j}^{\mathrm{obs}},\delta_{j},\varepsilon_{j},\gamma_{j}\mid\bm{z}_{j})$. Once observing the indicators values and fixing the entry dates, we have
$\mathrm{Pr}\bigl(T_{j}^{\mathrm{obs}}=t\mid\delta_{j}=1,\bm{z}_{j}\bigr)$ $=\frac{f_{\bm{\theta}}(t\mid\bm{z}_{j})}{F_{\bm{\theta}}(\tau_{j}\mid\bm{z}_{j}\bigr)}\,$, (11)
$\mathrm{Pr}\bigl(T_{j}^{\mathrm{obs}}=t\mid\varepsilon_{j}=1\bigr)$ $=\frac{\tilde{g}_{\bm{\psi}}(t)}{\tilde{G}_{\bm{\psi}}(\tau_{j})}\,.$
It follows that event times and loss-to-follow-up times are sampled from their respective truncated PDFs in (11), whereas interim censoring times are kept fixed at their realized values, thus ignoring how they are distirbuted. Sampling event times and loss-to-follow-up times from (11) can be carried out straightforwardly via inverse-transform. This approach is flexible and accommodates a wide range of lifetime distributions, requiring only evaluation of the corresponding truncated PDF and its inverse [9] [20].See Appendix, Section 7.5 for further details. In each bootstrap replicate, we obtain $\widehat{\bm{\theta}}$ and $\widehat{\bm{\psi}}$ by maximizing the likelihood in (10) under the same assumptions, and then compute $\pi_{j}$ at a future prediction horizon $\Delta t$ accordingly.
4 Simulation Study
We conducted two comprehensive simulation studies to globally assess the operating characteristics of the proposed method under various conditions of interest in the interim analysis contest. Our simulations not only assess the impact of the proposed extensions, but also investigate the behavior of the original DB-MC method.
4.1 Data generating process
We considered two simulation studies with different data generating processes and experimental conditions, namely $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$. In $\mathcal{S}_{1}$ data were generated from Weibull proportional hazards model with uniform entry dates over three years of recruitment period. In addition, we introduce an exponential loss-to-follow-up mechanism that is correlated with entry time, so that patients enrolled earlier are more likely to be loss to follow-up than those enrolled later. To induce this dependence while keeping the entry-time and loss-to-follow-up marginals PDFs parameterized separately, we use NORTA sampling (NORmal To Anything) [14] and calibrate the target correlation via a bisection procedure [5] with tolerance 0.001 and a maximum of 100 iterations with sample size 20000. The Weibull shape parameter was fixed at 0.6 (decreasing hazard). The baseline scale parameter was varied so that the proportion of interim censored observations matched the target censoring proportion under a specific value of the hazard ratio (see next Section), using also a bisection procedure [5] with starting value 0, tolerance 0.001, and a maximum of 300 iterations. We considered treatment as covariate, assuming a 1:1 randomization ratio generated from a Bernoulli distribution with parameter 0.5. The Poisson–binomial distribution was approximated in each bootstrap sample by a Poisson distribution, as Hong [39] showed this approximation to be highly accurate for a moderately large number of Bernoulli trials (e.g. >200) while being computationally more efficient than exact methods.
In $\mathcal{S}_{2}$, the data-generating process is similar $\mathcal{S}_{1}$ but we imposed only administrative censoring at the interim cutoff. Full details of the generating mechanisms and the resulting parameter values are reported in the Appendix, Section 7.6.
4.2 Factors in the simulations
In $\mathcal{S}_{1}$, the performance of the proposed method is evaluated over 72 scenarios defined by a combination of the following factors: the correlation between the loss to follow-up distribution and the entry dates $\rho=\{0.1,0.5\}$, the follow-up maturity at the time of prediction, $t_{c}=\{1,2,4\}$, expressed as the time from accrual end to the interim look; the ratio between the baseline Weibull scale and the loss to follow-up rate, $k=\{0.1,0.2,0.3\}$; the strength of the treatment effect, expressed by the hazard ratio, $HR=\{0.2,0.8\}$; and the future prediction horizon, $\Delta t=\{1,4\}$. The proportion of interim censored observations is fixed at 0.5. Here, we evaluate prediction intervals under correct model specification.
In $\mathcal{S}_{2}$, the methodology is evaluated over 24 scenarios defined by a combination of the following factors: the proportion of interim censored observations, $0.2=\{0.2,0.8\}$ and $t_{c}$, $HR$, $\Delta t$ using the same values defined in $\mathcal{S}_{1}$. We evaluated prediction intervals obtained from the Weibull model, Royston–Parmar models with 1, 2, and 3 knots placed at the 25th, 50th, and 75th percentiles of the uncensored event times and the Generalized Gamma model [67], which encompasses Weibull and log-normal as special cases [59] [19]. Following Mori et al. [56], we decided to include Royston–Parmar models [62] as a flexible extension of standard parametric regression models. These models generalize the linear relationship between a transformed survival function, i.e. a link function $g(S(t))$ and log time $y=\log(t)$ on a chosen covariates effect scale (e.g. proportional hazard scale) by modeling this relationship with restricted cubic splines. In this sense, they extend the Weibull proportional hazards model, the log-logistic proportional odds model, and the log-normal linear probit model. Despite their added flexibility, Royston–Parmar models preserve an explicit baseline specification, allowing extrapolation beyond the last observed event. Flexibility is controlled by the number of internal knots placed between boundary knots [62].
For each combination of factors, the metrics were computed and averaged on $N=1000$ Monte Carlo samples with sample size $n=1000$. For each simulated dataset, 100 bootstrap resamples were used to construct the prediction intervals. Although 100 bootstrap replicates is relatively modest, this choice was necessary to keep the overall computational burden manageable across all scenarios.
4.3 Performance metrics
The prediction intervals obtained within the DB-MC framework enjoy the classical frequentist interpretation: they correspond to the long‐run proportion of times that $Y$ falls in a random region $\mathrm{PI}_{1-\alpha}(\mathscr{D})$, i.e.the prediction interval, under repeated sampling from the same data‐generating mechanism [8, 50]. Coverage probability (CP) is a useful performance metric to evaluate an interval estimator. Since $Y$ is also random, there are two natural notions of CP associated with a prediction interval procedure [10, 54]. First, for the observed data $\boldsymbol{D}$ there is the coverage probability of the realized prediction interval, i.e., the conditional coverage probability (CCP)
$CCP=P\bigl(Y\in\mathrm{PI}_{1-\alpha}(\bm{D})\mid\varphi(\boldsymbol{\vartheta})\bigr).$ (12)
[…]
$UCP$ $=\mathrm{E}_{\mathscr{D}}\bigl[CCP\mid\varphi(\boldsymbol{\vartheta})\bigr]$ (13)
$=\mathrm{E}_{\mathscr{D}}\!\left[\Pr\!\left\{\,Y\in\mathrm{PI}_{1-\alpha}(\boldsymbol{D})\,\middle|\,\varphi(\boldsymbol{\vartheta})\right\}\right]$
$=P\bigl(Y\in\mathrm{PI}_{1-\alpha}(\mathscr{D})\mid\varphi(\boldsymbol{\vartheta})\bigr).$
When a closed‐form expression for (13) is not available, UCP can be approximated by Monte Carlo simulation [53] [55]) as
$UCP\approx\frac{1}{N}\sum_{k=1}^{N}P(Y\in\mathrm{PI}_{1-\alpha}(\boldsymbol{D}_{k}^{*})),$ (14)
where $\bm{D}^{*}_{1},\ldots\bm{D}^{*}_{N}$ are simulated data from a given model. In our simulation studies, $\bm{\vartheta}=(\bm{\theta},\bm{\psi},\bm{\phi})$ and $\varphi(\boldsymbol{\vartheta})={\bm{\pi}}$. The UCP of the prediction intervals is estimated under the factors described in Section 4.2 using the following steps:
  1. 1.
    Generate $N$ independent samples $\boldsymbol{D}_{k}^{*}$ under the data‐generating processes detailed in Section 4.1.
  2. 2.
    For each simulated sample, obtain the prediction interval $\mathrm{PI}_{1-\alpha}(\boldsymbol{D}^{*}_{k})$ by applying the proposed parametric bootstrap scheme to obtain the conditional CMF of $Y$ and by taking the corresponding $\alpha/2$ and $1-\alpha/2$ quantiles. For $\mathcal{S}_{2}$ we obtain the quantiles under different parametric models (See Section 4.1).
  3. 3.
    For each $\boldsymbol{D}^{*}_{k}$, we evaluate the CCP using the conditional distribution of $Y$ evaluated at the true $\bm{\pi}$}, i.e. the $\bm{\pi}$} evaluated on the simulated dataset $\boldsymbol{D}^{*}_{k}$ at the design value of $\bm{\vartheta}$.
  4. 4.
    Average these evaluations as in (14).
The closer the estimated prediction bounds are to the $\alpha/2$ and $(1-\alpha/2)$ quantiles of the conditional distribution of $Y$ evaluated at the true $\bm{\pi}$ in each sample, the better the resulting coverage. For this reason, we additionally assess the relative bias of the estimated lower and upper prediction bounds, as well as the ratio between the width of the estimated interval and that of the true interval for the simulated dataset. These measures are respectively defined as
$\begin{gathered}\frac{y_{\alpha}\bigl(\boldsymbol{D}_{k}^{*}\bigr)-y_{\alpha}\bigl(\boldsymbol{D}_{k}^{*},{\bm{\pi}})}{y_{\alpha}\bigl(\boldsymbol{D}_{k}^{*},{\bm{\pi}}\bigr)}\,,\\ \\[4.0pt] \frac{\mathrm{PI}_{1-\alpha}(\boldsymbol{D}_{k}^{*})}{\mathrm{PI}_{1-\alpha}(\boldsymbol{D}_{k}^{*},{\bm{\pi}})}\,.\end{gathered}$ (15)
where $y_{\alpha}\bigl(\boldsymbol{D}_{k}^{*},\bm{\pi})$ and $\mathrm{PI}_{1-\alpha}(\boldsymbol{D}_{k}^{*},{\bm{\pi}})$ are the $\alpha$ quantile and the prediction interval obtained on the simulated dataset $\bm{D}_{k}^{*}$ at the true value of $\bm{\pi}$}.
4.3.1 Software
All simulations were performed in R 4.5.1 [61] and analyzed with the rsimsum package [27]. Simulations from an (uncensored) Weibull regression model under a proportional hazards specification were generated using the simsurv package [13]. Weibull models were fitted with the survival package [70], whereas Generalized Gamma and Royston–Parmar models under proportional hazards were fitted with flexsurv [41]. Random variates from parametric survival models were obtained via the quantile functions implemented in flexsurv. Numerical integration was performed using the R function integrate(), which perform adaptive quadrature [61].
For the case study, nonparametric and semiparametric survival analyses were conducted using survival and muhaz [35]. Weibull, log-normal, and log-logistic regression models were fitted with survival, while Generalized Gamma and Royston–Parmar models were fitted with flexsurv. Evaluation of the Poisson–Binomial distribution was carried out using the poibin package [38].
4.4 Results
In $\mathcal{S}_{1}$, overall coverage is close to the nominal level and it varies mainly with the prediction horizon $\Delta t$ and follow-up maturity $t_{c}$ (Figure 1). Longer horizons are associated with lower coverage, whereas more mature follow-up yields a modest improvement. Loss to follow-up also slightly reduces coverage, with a stronger impact for longer-term predictions. This pattern reflects a trade-off in information: higher loss to follow-up improves estimation of the censoring process but reduces information for estimating the event-time distribution.
Prediction intervals widen as $\Delta t$ increases and $t_{c}$ decreases, particularly when loss to follow-up is modest. However, wider intervals do not necessarily translate into higher coverage. Worsening occur mainly when both interval bounds shift in the same direction, producing intervals that are systematically biased upward (i.e., overestimating future event counts). Note that interval coverage here also reflects which portion of the support of the conditional predictive distribution is captured by the interval.
In $\mathcal{S}_{2}$ under correct model specification, the prediction intervals maintained coverage close to the nominal level across almost all scenario combinations (Figure 2). Coverage deteriorated slightly only when follow-up maturity was low, the treatment effect was modest - hazard ratio equal to 0.8 - and predictions were made far beyond the interim time, with a maximal loss of about 5 percentage points.
Over-parameterised survival models were generally not inferior in terms of coverage compared with their nested counterparts, but this was achieved at the cost of wider intervals and systematically downward-biased bounds particularly in settings with substantial interim censoring. In contrast, the upper bounds of the predictive intervals behaved very similarly across models. Notably, for over-parameterised models the coverage deteriorated in scenarios with low follow-up maturity, especially when administrative censoring was limited, the hazard ratio was modest, and predictions were made far into the future. This loss of coverage was driven by a systematic underestimation of both lower and upper predictive limits, so that the resulting intervals tended to concentrate on the support of the conditional predictive distribution with relatively low true probability.
These patterns were particularly evident for the Royston–Parmar models (Figure 6, Appendix): increasing the number of spline knots, and hence model complexity, did not systematically improve performance. As follow-up maturity increased, coverage improved, but at the expense of wider intervals compared with the Generalized Gamma model.
Figure 1: Study $\mathcal{S}_{1}$. Performance metrics of the 95% two-sided prediction intervals under different factors. $\Delta t$ denotes the prediction horizon, $t_{c}$ the time of interim from accrual end, $\rho$ the correlation between loss to follow-up times and entry dates, HRHR is the hazard ratio and $k$ the ratio between the baseline scale parameter and the loss to follow-up rate.
Figure 2: Study $\mathcal{S}_{2}$. Performance metrics of the 95% two-sided prediction intervals under different models and administrative censoring only at $t_{c}$. $\Delta t$ denotes the prediction horizon, $p(t_{c})$ denotes the proportion of interim censored patients, $t_{c}$ the time of interim from accrual end, HRHR is the hazard ratio and $k$ the ratio between the baseline event rate and the loss to follow-up rate.
5 Case study in Pediatric ALL
To illustrate the proposed method, we considered data from a multicenter phase III trial on pediatric ALL published by Conter et al. [17]. The AIEOP-BFM ALL 2009 protocol was an international, open-label, randomized phase III study conducted across seven countries and coordinated by the AIEOP-BFM consortium. The trial evaluated whether intensifying asparaginase exposure during early consolidation could reduce minimal residual disease (MRD) and improve outcomes in children and adolescents with high-risk ALL. In the experimental arm, patients received four additional weekly doses of pegylated L-asparaginase during Consolidation IB on top of standard chemotherapy, while the control arm followed the same regimen without PEG-ASNase. Between June 2010 and February 2017, 6136 children (aged 1–17 years) with Philadelphia-negative ALL were enrolled; 1097 met high-risk criteria at the end of induction and 809 were randomized, 404 experimental, 405 control. The primary endpoint was the proportion of patients with PCR-MRD $\geq 5$x $10^{-4}$ at the end of consolidation. Secondary endpoints included event-free survival (EFS), overall survival (OS), cumulative incidence of relapse, and toxicity.
For the application of our method, two-sided prediction intervals at the 0.95 confidence level were obtained for the the number of additional EFS events from June 1, 2017 (accrual closure) and updated every six months until June 1, 2021, when the final analysis was scheduled. The full dataset was used to evaluate if the observed events are included by the prediction intervals. EFS is defined as the time from random assignment to the date of the first event (resistance to treatment, relapse, death, or second malignant neoplasm, whichever occurred first) or to the last follow-up. Since no loss to follow-up occurred at the time of the final analysis, all patients who had not experienced an EFS event by June 1, 2017 were considered administratively censored at that date. The presence of interim censoring only leads to a simplified formulation of (8) and the sampling scheme, as detailed in the Appendix, Section 7.3. On June 1 2017, 164 events had occurred among 809 enrolled patients. By the final analysis, 62 further events had been observed, for a total of 226 events over the entire trial.
[…]
Figure 3 shows the relationship between different transformations of the survival function and natural logarithm of time: straight parallel lines suggest correct specification of both baseline survival and covariates effect scale. These plots showed that there is no linear relationships between the transformed survival and the logarithm of time, suggesting that Royston-Parmar models may be adequate. In addition, PH assumptions are untenable. The smoothed hazard plot (Figure 8, Appendix) indicates decreasing hazards converging later in the follow-up.
Prediction intervals for the survival function obtained from simple parametric regression models (Weibull, log-logistic, and log-normal) were broadly similar across the prediction window and were consistently shifted upwards relative to the flexible counterparts (Figure 4). Within the Royston–Parmar models, different link functions (PH, PO, LP) produced comparable intervals and displayed a coherent pattern as the number of internal knots increased: specifications with 2–3 knots yielded wider and more downward-shifting prediction intervals as the prediction horizon lengthened, compared with the 1-knot models.
Model-fit comparisons based on information criteria (Table 1) highlight the trade-off between parsimony and goodness of fit. Although the maximized log-likelihoods were very similar across models, BIC tended to favor more parsimonious specifications, whereas AIC mainly separated simple parametric models from their more flexible counterparts. For instance, the log-normal model had the third-lowest BIC but one of the third-highest AIC values.
Model discrimination should also consider extrapolative performance. Inspection of the fitted hazard functions (Figures 9 to Figures 15, Appendix) showed that models imply different hazards, yet these differences matter for survival extrapolation primarily through the cumulative hazard function (Figures 16 to Figures 18, Appendix). Moreover, despite distinct baseline hazard shapes and different covariates effect scales, Weibull, log-logistic, and log-normal models have broadly similar standardized distributions within the log-location-scale familiy i.e., Gumbell, Logistic, Normal. Thus, they are expected to exhibit similar extrapolative performance despite having different fit. Analogous considerations apply to Royston–Parmar specifications.
Figure 3: Different survival function transformations vs. log-time by treatment arm, 1 = Active treatment, 0 = standard of care. Panel A shows the log cumulative hazard of EFS $\mathrm{log}(-\mathrm{log}(S(t)))$; Panel B shows the cumulative probit of EFS $\Phi^{-1}(S(t))$, where $\Phi^{-1}$ is CDF of the standard normal distribution; Panel C shows the cumulative logit of EFS $\mathrm{log}(1-S(t)/S(t))$.
Figure 4: Two-sided 95% prediction intervals for the additional number of EFS events from June 1st 2017 to June 1st 2021 . Black diamonds indicate the observed counts. Panel A shows intervals under the Weibull model and its flexible specifications; Panel B shows under the log-normal model and its flexible specifications; Panel C shows under the log-logistic model and its flexible specifications. RP = Royston–Parmar; $k$ denotes the number of knots.
Table 1: Models fit measures
Model $q$ logLik AIC BIC
Generalized Gamma 4 562.9134-562.9134 1133.827 1152.585
RP – PO(1) 4 564.2135-564.2135 1136.427 1155.185
Log-normal 3 567.7988-567.7988 1141.598 1155.666
RP – PH(1) 4 564.6479-564.6479 1137.296 1156.054
RP – LP(1) 4 564.6479-564.6479 1137.296 1156.054
RP – PH(2) 5 562.1776-562.1776 1134.355 1157.803
RP – LP(2) 5 562.1776-562.1776 1134.355 1157.803
RP – PO(2) 5 563.1182-563.1182 1136.236 1159.684
RP – PH(3) 6 561.4914-561.4914 1134.983 1163.120
RP – LP(3) 6 561.4914-561.4914 1134.983 1163.120
RP – PO(3) 6 562.4590-562.4590 1136.918 1165.056
Log-logistic 3 573.9548-573.9548 1153.910 1167.978
Weibull 3 576.7842-576.7842 1159.568 1173.637
Note: Models are ordered by increasing BIC. $q$ denotes the number of parameters. RP = Royston–Parmar model, PH($k$) = Proportional Hazards, PO($k$) = Proportional Odds, LP($k$) = Linear Probit, where $k$ denotes the number of knots.
6 Discussion and Conclusions
In this work, we propose a general framework to predict the additional number of events during interim monitoring of clinical trials with time-to-event endpoints, together with corresponding prediction intervals. The framework accommodates patient-level covariates, loss to follow-up, and possible dependence between entry times and loss to follow-up. We adapt and extend to the interim analysis setting the DB–MC approach of Tian et al. [71], originally developed in reliability engineering to construct prediction intervals for the number of future item failures. Prediction intervals are obtained as the quantiles of a parametric-bootstrap estimator of the conditional CMF of the future event count. Importantly, the parametric bootstrap is implemented conditional on the observed event and censoring indicators, thereby preserving the information available at the interim analysis.
The Poisson–binomial representation of the future event count provides a flexible way to model heterogeneous, subject-specific conditional probabilities of experiencing the event within the future time window, driven by staggered entry and, more generally, by covariate profiles. This formulation also naturally allows different parametric models for event times and censoring times. Embedding such models within the sampling scheme proposed is straightforward since it requires simulating event and loss to follow-up times from the chosen truncated parametric distributions. This yields substantial flexibility, leveraging a broad class of parametric survival models available in the literature [21, 51]. In this work we assumed a common dropout mechanism across covariate profiles, but the approach readily extends to settings where dropout also varies by covariates (e.g., treatment arm). The key requirement is that, conditional on the covariates, the event time and the loss-to-follow-up time are independent and their distributions share no parameters.
We examined the operating characteristics of the proposed method across a range of factors of interested in interim monitoring settings and with a case study on pediatric ALL. Follow-up maturity and prediction horizon emerged as key determinants of interval performance: more mature follow-up and shorter horizons led to better coverage, whereas longer horizons reduced coverage. Loss to follow-up produced modest reduction in coverage, which became more pronounced for longer prediction horizons. This reflects an information trade-off: while additional loss to follow-up can inform the censoring model, it simultaneously reduces the information for estimating the event-time distribution because fewer events are observed.
Across simulations, using more flexible parametric extensions of simpler models did not systematically improve prediction performance. This result aligns with those of Mori et al. [56]. Although a more general model could be a safe choice if assumptions on the survival distribution are difficult to ascertain, the risk of overfitting should be considered in settings with low follow-up maturity, even when the proportion censored is small. The analysis of the real case suggests that model choice should be guided not only by information criteria but also by extrapolative behavior beyond the maximum follow-up observed at interim assessed via graphical inspection. Models implying different survival and hazard shapes may exhibit similar extrapolation, e.g., members of the log-location–scale family. The ability of the proposed framework to accommodate multiple parametric specifications also facilitates sensitivity analyses for predicted event counts under different distributions.
A further remark concerns the inferential perspective adopted in this work. In this work, inference is conditional on the realized interim dataset. In a frequentist setting, uncertainty is quantified by viewing the observed data as one realization from a hypothetical sequence of replications conducted under the same conditions. Here the same principle is applied to a future state of the same trial: the target is the distribution of what will be observed at a later calendar time, i.e., a within-sample prediction problem. This implicitly requires that the information available at the interim time is representative of what will be observed at later looks. The proposed bootstrap scheme is designed to mirror this perspective.
Further research could address the limitations underlying our approach and connect it to a set of active areas of research. First, we assume that the data at the interim analysis is free of reporting delays. In multicenter trials, however, delays in data entry and event reporting are common and can materially affect interim monitoring and prediction [73, 75, 4]. Extensions that explicitly account for delayed reporting would therefore be valuable. Second, we focus on settings in which accrual is complete at the prediction time. In many trials, interim analyses occur while recruitment is ongoing, so a practically relevant extension is joint prediction of both future accrual and future events as often considered in literature on time to endpoint maturation. Third, we assume access to unblinded, patient-level information. In practice, interim analyses are frequently conducted under blinding constraints to preserve trial integrity, limiting the availability of individual-level covariates or even treatment indicators. Related work on time to endpoint maturation prediction under blinded data [22, 26, 84] suggests promising directions to adapt our framework to blinded settings. Regarding the survival model, we restrict attention to regression-based parametric specifications without explicitly modeling potential violations of the covariates effect scale (e.g., departures from proportional hazards, proportional odds, or accelerated failure time assumptions). While in the motivating application ignoring such violations did not lead to marked differences in prediction intervals, the generalizability of this observation warrants further investigation. Lastly, our work may benefit from the extensive literature on parametric survival extrapolation from clinical trials for economic evaluation [49, 64, 29, 18, 48, 68, 40, 58, 15], where model choice and extrapolative performance are central concerns. Leveraging insights from this literature could be a promising avenue for future work.